OpenStudy (anonymous):

why does x/2-x equal (-x)/2

OpenStudy (anonymous):

x/2 - x = (x/2) - (2x/2) = (-x)/2

OpenStudy (anonymous):

where did you get 2x/2

OpenStudy (anonymous):

since you are subtracting fractions you have to have a common denominator, whic in this case is 2, so x becomes 2x/2

OpenStudy (anonymous):

ohhh thanks so much

OpenStudy (anonymous):

You're welcome